The height of a cylindrical tank, containing 20 lb of gas at 70 psia and 115 Farenheit, IS TRIPLE ITS DIAMETER; the gas constant is R = 45 ft lbf/lbm R. find.


1
Expert's answer
2021-12-20T10:28:42-0500

Solution:


"V=\frac{mRT}p,"

"V=Sh=\frac{\pi d^2}4 h=\frac{\pi h^3}{36},"

"h=\sqrt[3]{\frac{36 mRT}{\pi p}}=8.02~ft."